OpenStudy (shubhamsrg):

Time pass ques: Use 5 zeroes with any mathematical functions to arrive at 14. PS: This has a lame solution. -_-

OpenStudy (anonymous):

f(x) = x - 100000 f(10014) = 14

OpenStudy (shubhamsrg):

cool.. B|

OpenStudy (shubhamsrg):

But NO :|

Parth (parthkohli):

Let \(f(n) = n + 14\). Then \(f(0 + 0 + 0 + 0 + 0) = 14\)

Parth (parthkohli):

B-)

Parth (parthkohli):

My solution is lamer!

OpenStudy (shubhamsrg):

-_-

OpenStudy (mathstudent55):

f(x) = 1400000x f(0.00001) = 14 I used 5 zeros twice :-)

OpenStudy (shubhamsrg):

good for you :)

OpenStudy (anonymous):

S(S(S(S(S(S(S(S(S(S(S(S(S(S(0+0+0+0+0)))))))))))))) This one is the lamest :D

OpenStudy (shubhamsrg):

what is this anyways ?

OpenStudy (anonymous):

or \(p_{0!}p_{0!+0!+0!+0!}\), where \(p_n\) denotes the nth prime?

OpenStudy (anonymous):

S(n) = n + 1, succesor function

OpenStudy (shubhamsrg):

aha..hmm I won;t say thats more lame, I'd say thats equally lame relative to my soln.! -_-

OpenStudy (shubhamsrg):

but still, its cool.. B|

OpenStudy (anonymous):

What's your solution? :p

OpenStudy (shubhamsrg):

I'll message you.

Parth (parthkohli):

So the answer is...\[\left\lfloor \tan\left(\dfrac{0! + 0! + 0!}{0! + 0!}\right)\right\rfloor\]